OpenStudy (anonymous):

Use a series to express the following number as a ratio of integers: 0.73737373

OpenStudy (anonymous):

\[\frac{73}{100}+\frac{73}{10000}+\frac{73}{1000000}+...\]

OpenStudy (jlastino):

Can also be 73/99 If you're doing infinite geometric series

OpenStudy (anonymous):

you get \[\frac{73}{100}\sum_{k=0}^{\infty}(\frac{1}{100})^k\]

OpenStudy (anonymous):

and as jlastino wrote, if you add this up you get \(\frac{73}{99}\)

OpenStudy (anonymous):

Thank you guys! Can I write 0.2222 as: \[\sum_{1}^{\infty} 2 (1/10)^{n} = 2/10 \div 1-1/10\] ?? Oh and how are you guys making your fractions look like that?

OpenStudy (jlastino):

There's a shortcut for writing repeating decimals haha You take the repeating part then divide it by a # of 9's (The number depends on how many numbers repeat) Which in this case 2 So It can be written as 2/9

OpenStudy (anonymous):

Oooooo. My teacher may have said that. But I missed this lesson in class, so Im teaching myself. Thanks again. I SHALL post more!

OpenStudy (anonymous):

fractions you write via \frac{a}{b}

OpenStudy (anonymous):

and no agony for repeating decimals. \[\overline{.1}=\frac{1}{9}\] \[\overline{.2}=\frac{2}{9}\] \[\overline{.3}=\frac{3}{9}\] \[\overline{.4}=\frac{4}{9}\] \[\overline{.5}=\frac{5}{9}\] etc

OpenStudy (anonymous):

Lol, this is great!

OpenStudy (jlastino):

More examples: 0.123123123 = 123/999 0.14331433= 1433/9999 0.515151 = 51/99

OpenStudy (anonymous):

What about 6.254, with "54" repeating? I have: \[62\sum_{n=1}^{\infty} 54 (\frac{1}{1000})^n\] Now i just gotta figure out how to solve.

OpenStudy (jlastino):

I forgot how to solve (sadly) I just do the shortcut T_T It'll be 62/10 + 54/990 Ask satelite for help on the solution

OpenStudy (anonymous):

I have the same answer: 62/10 + 54/999= 6.2540